# Observing radio transients with Phased ALMA: Pulses from the Galactic Centre magnetar

### v2

@@ 4.2 Energy distribution
We construct the cumulative energy distribution by sorting the events and then counting the number above an energy threshold, until all the pulses are contained. We include only events above the completeness limit 777This is the energy limit of a burst having the minimum detectable fluence given a telescope sensitivity., which is 1120 mJy ms using a 51 K and = 1.15 K Jy-1 (Liu et al., 2021) and considering the pulse width as the mean FWHM of our pulses, 0.07 s. We fit a power-law to the cumulative energy distribution using least squares minimisation considering that the event rate scales with the energy as . The result is a power-law exponent of .
@@ 5 Discussion
Our sample adds five pulses in addition the three previously reported by Liu et al. (2021), mainly due to the difference in the candidate selection criteria. In our work, we defined the SNR threshold as 6, based on an statistical analysis of the noise floor. For each of these candidates the four Stokes parameter plus linear and circular polarisation were inspected. Polarisation information was crucial as the dispersion delay between frequency channel was negligible at ALMA Band-3 frequencies.
@@
Among the pulses, five of them: P1, P2, P3, P5 and P8, exhibit nearly 100 of linearly polarised flux, while the remaining (P4, P6 and P7) have values close to 50. Also, for the majority of them there is a degree of circular polarisation surrounding 30. Regarding the pulse morphology, as seen in Fig. 1, most of the pulses have at least two components, which vary in relative intensity. All pulses show emission across the entire 2 GHz bandwidth. Broadband emission is one of the features of radio pulsars (Lorimer & Kramer, 2004).
@@
This is the first time the energy properties of the pulses from the GC magnetar have been explored at high frequencies. Previous studies were conducted at 3.1 GHz (Yan et al., 2018) and 8.6 GHz (Yan et al., 2015). Common fits to the energy distribution of single pulses are log-normal and power-law distributions. For the power-law, which is the method used in our study, the defining parameter is , the power-law exponent such that , with the rate of events above a given energy value . This definition corresponds to the cumulative energy distribution, which is broadly used in cases where the number of pulses is limited, such as the case of FRBs.
@@
As shown in Fig. 3, our sample is well fit by = -2.4 0.1. For the same source, at centimetric wavelengths, the energy distribution of single pulses was found to be well fitted by a log-normal distribution with = 1.34 and = 0.57 at 3.1 GHz (Yan et al., 2015), and = 0.72 and = 0.22 at 8.6 GHz (Yan et al., 2018). The difference with our result is expected given the sample size, where the authors have considered 1900+ and 2300+ pulses, respectively. However, it is known that pulsars show a bimodal behaviour: their regular emission is well modelled with a log-normal distribution (Burke-Spolaor et al., 2012), while giant pulses are modelled by a power-law distribution (Karuppusamy et al., 2010). In the case of the GC magnetar, previous studies on the single pulse peak flux densities of giant pulses at 8.4 GHz, found that they are not well fitted by a log-normal distribution (Pearlman et al., 2018). Given the steep spectral index in pulsars, the observable events at millimetre wavelengths could correspond to the most energetic pulses. Alternatively, a power-law relation for the cumulative energy distribution would be the result of the limited energy range of our pulses. Extending the observations with on-source time and frequency coverage will increase the sample and permit us to have a better general picture of the behaviour of PSR 1745-2900.
@@
Our is consistent with other magnetars’ energy distributions at cm-wavelengths. For XTE J1810197 (PSR J18091943) giant pulses, it has been reported values between -2.1 and -7.68 at 1.4 GHz, between -1.78 and -2.86 at 4.9 GHz, and between -0.89 and -1.89 at 8.35 GHz. The differences could be attributed to its intrinsic magnetar variability and the observations at different frequencies (see Table 2 in Serylak et al. 2009).
@@
Importantly, our results fill a critical gap by characterizing magnetar radio emission across a scarcely explored frequency window, offering valuable insights into potential changes in their emission mechanisms. Magnetar radio emission –known to be coherent at cm-wavelengths– persists into the mm-band, as demonstrated by our work, Liu et al. (2021), and Torne et al. (2017). However, it remains unclear exactly where a transition to an incoherent emission occurs. This transition is expected to take place in the mm to sub‑mm range, yet current time-domain observational facilities are few. In this context, PAM, given its sensitivity, holds significant potential.
@@
In regard to FRBs, there are three sources with seemly periodic activity windows: FRB 20121102A with a periodicity of 159.3 days (Braga et al., 2025), FRB 20180916B with 16.35 days (Chime/Frb Collaboration et al., 2020) and recently FRB 20240209A with 126 days (Pal, 2025). The periodicity corresponds to the cycle of the alternating on-and-off activity. Naturally, observations carried out at their peak activity are more likely to result in a detection. FRB 20121102A has values between (Cruces et al., 2021) and (Gourdji et al., 2019). However, Li et al. (2021) shows using a large sample of 1600+ bursts detected by the Five-hundred-metre Aperture Spherical Telescope (FAST), that the energy distribution is bimodal and best fit by a combination of a log-normal and a Cauchy function. In that work, the analysis was done over the histogram of the differential energy distribution rather than the cumulative distribution, given the large sample of bursts available. In Li et al. (2021), when analysing the high energy component, it was found that the sub-sample is well fit by a power-law with . It is important to note that in our work we reported a limited sample of 8 pulses. With an even larger sample size of hundreds or more pulses the differential energy distribution could be as well be computed.
@@
For another active repeater, FRB 20201124A, it has beeen reported -values such as -1.5 0.1 in Xu et al. (2022) and -1.95 0.01 to -2.25 0.02 in Kirsten et al. (2024), both at cm-wavelengths. A steeper value of is often seen at higher energies in repeating FRBs (Kirsten et al., 2024), reaching values as steep as in the high-energy tail (Kumar et al., 2023). Recent studies of FRBs’ energy distribution have reported a value of -1.96 0.15 for the pure power-law (Arcus et al., 2024) using FRB data between 0.83 and 1.6 GHz.
@@
The typical isotropic energies of FRBs are in the range of erg (CHIME/FRB Collaboration et al., 2020). Comparing the energy of our pulses, we find a gap of at least seven orders of magnitude. Nevertheless, the hypothesis of FRBs being extragalactic magnetars (Sobacchi et al., 2024) is supported by the activity of the Galactic magnetar SGR 1935+2154, which emitted events with energies in the range of 1034 (Margalit et al., 2019; Bochenek et al., 2020). This values are closer to the ones seen for FRBs, but still at least two orders of magnitude below. There are some phenomena that could explain the energy gap. One is multiple magnifications of the pulses due to plasma lensing, especially in environments where magnetars are orbiting a compact source such as a massive black hole (Pen & Connor, 2015; Pearlman et al., 2018). In this regard, a promising laboratory to test this scenario is the GC. While for most radio pulsars and FRBs the monitoring is done at lower frequency bands, complex regions such as the GC are hard to reach with centimetric waves as the pulses are heavily affected by propagation effects such as scattering and dispersion. Therefore higher frequencies allow us to probe deeper into these complex regions. The natural disadvantage is the lower luminosities expected at higher frequencies. Our study is specially remarkable given the detection of several single pulses from the GC magnetar at 86 GHz, which allow us to characterise the source, explore the potential of ALMA for radio transients and also to stablish a direct link to what we could expect for FRBs at higher frequencies, if they are indeed originated by magnetars.
@@
Although most pulsars display a steep spectral index, magnetars such as PSR 1745-2900, have a flat or inverted one (Torne et al., 2017; Pearlman et al., 2018). Their spectral index could also change in time (see e.g. Champion et al. 2020). Torne et al. (2015) reports a value of for the GC magnetar, that changes to in a different epoch (Torne et al., 2017). This positive value could reflect a turn up, related with the transition from coherent to incoherent radio emission. Observing with high sensitivity instruments such as ALMA could be helpful to study the gap between centimetre and millimetre radio wavelengths, particularly with its new Band-1 setup.
5.1 PAM Band-1 predictions
@@
Based on our detections, we report a Poissonian rate of = 4 events/hr for PSR J1745-2900 at ALMA Band-3. We can estimate the expected rate at Band-1 (42 GHz) using:
@@
Here corresponds to the expected event rate for a given fluence threshold at Band-X, respectively. The fluence completeness limit using 37 antennas of 12m at Band-3 is = 1330 mJy ms and at Band-1 = 1200 mJy ms, considering a pulse width of 0.1 s. The exponent is the power-law value derived from the cumulative energy distribution discussed in Sec. 4.2. We obtain a rate of = 5 events/hr. However, if we instead consider a total of 49 antennas for the phasing, then the expected rate rises to = 10 events/hr, as the fluence threshold drops to = 900 mJy ms. Certainly, having a larger sample of pulses will provide more constraining values for their energy and frequency behaviour. In particular, measuring the spectral index of the source at different frequencies will tell us whether there is a turn-up point when we transition to incoherent emission mechanisms.
@@ 5.1.1 FRBs
If we consider the obtained for the GC magnetar and apply it to some actively repeating FRBs, we can estimate the event rates at ALMA bands. In this case, we re-calculated the fluence threshold for ALMA assuming bursts of 1 ms duration (typical for FRBs), and obtain = 90 mJy ms 888Using a maximum of 49 antennas of the same 12m size.. We take FRB 20121102A, whose reported rates at L-band are 122 events/hr (Li et al., 2021, 2022) and 21816 events/hr (Jahns et al., 2023). This leads to a rate between 1 and 3 events/hr, respectively. If on the contrary, we take the flatter value of reported for FRB 20121102A at L-band in Cruces et al. (2021), we estimate instead between 10 and 32 events/hr. Noteworthy is that this rates consider observations at the peak activity of the source.
We highlight that PAM Band-1 could serve as a bridge to millimetre wavelengths, having profound implications in constraining the FRBs’ progenitor source. If detections are achieved, it will also be key to test the proposed frequency dependence of the activity window (chromaticity) of periodic FRBs. As discussed in Bethapudi et al. (2023), the activity window seems to starts earlier and be narrower at higher frequencies. As shown in Fig. 4, as of now, ALMA could potentially detect 19 repeating FRBs. This number is expected to increase rapidly given the new surveying instruments in the southern hemisphere such as MeerKAT (Jonas & MeerKAT Team, 2016).
@@ 5.1.2 Pulsars
We can also estimate the number of radio pulsars detectable in integrated pulse profile mode (also known as folding mode), which is obtained by adding hundreds to thousands of pulses using the spin period and other parameters listed in their ephemeris (see e.g. Table 3). We derived the minimum sensitivity for Band-3 using the radiometer equation for pulsars and considering telescope parameters such as = 74 K, = 1.08 K/Jy (Cortes et al., 2024) for Band-3, a phasing efficiency of (Matthews et al., 2018), a minimum detected SNR of 6 and 1 hour of on-source integration time. For a pulsar spin period of 1 s with a duty cycle of 10%, we estimate that Band-3 system is sensitive to 0.060 mJy. If instead we use the parameters reported on Liu et al. (2021), we obtain 0.023 mJy. Assuming that Band-1’s system behaves similarly ( = 45 K, = 1.12 K/Jy in Cortes et al. 2024), then we expect a minimum detectable flux of 0.035 mJy, using the same 37 antenna setup. When increasing to 49 antennas, the minimum flux is 0.026 mJy.
In Fig. 5 we show the sources expected to be detectable by PAM Band-1. To this end, we selected all the known pulsars from the ATNF catalogue (Manchester et al., 2005) having flux density reported at 1.4 GHz or 2 GHz. We limited the sources detectable by ALMA based on a declination below 47° (Cortes et al., 2024). This reduced the list to 2301 pulsars.
We used the spectral index to estimate the flux of each source at 42 GHz (Band-1). To extrapolate we used a simple power-law following:
where corresponds to the flux at a given frequency in GHz, is the spectral index and is the frequency. For the sources which do not have reported spectral index, we draw a spectral index from a normal distribution, considering a mean of -1.6 and standard deviation equal to 0.54 (Jankowski et al., 2017b). We assumed a pulse width equal to 10 of the period and considered on-source integration times of 1, 1.5, and 2.5 hours. This would lead to the detection of 166 pulsars (7.2 of those visible to ALMA), 188 (8.2), and 209 (9.1), respectively. The sources detectable by PAM are the ones lying above the mentioned sensitivity lines in Fig. 5. Some remarkable detectable sources have been highlighted, such as the Vela pulsar, and three radio-loud magnetars, including PSR J17452900. It is important to mention that these surveys are unlikely to be sensitive to millisecond pulsars (MSPs, Liu et al. 2021).
@@ 6 Conclusions
We reprocessed archival data of APS at Band-3 from Sgr A∗. We detected eight pulses from the GC magnetar PSR J17452900 and studied their cumulative energy distribution with the objective to link this with FRBs. The complex environment where the GC magnetar is located serves us to test promising progenitor scenarios for FRBs, given its high electron density and turbulent magneto ionic properties.
We obtained an exponent of = 2.4 0.1 for the cumulative energy distribution of the GC magnetar. Using this exponent we predict a rate of 5 events/hr if we were to observe in Band-1, using 37 antennas of 12m. This rate increases to 10 events/hr with 49-12m antennas. For the periodic FRB 20121102A, at peak activity, we obtained a rate between 10 and 32 events/hr with PAM Band-1 using 49 antennas of 12m. Regarding pulsar detection, we estimated that there are roughly 166 pulsars potentially detectable within 1 hr integration. This number increases with more integration time.
Currently, there are 19 repeating FRBs which are visible to ALMA. Given its high sensitivity and time resolution, instruments like PAM are crucial to explore FRB’s frequency extend to test their progenitor scenarios. Importantly, high-frequency searches are not affected by propagation effects, such as scattering, which could hinder a detection at cm-wavelengths. It is expected that the Band-1 receiver, including polarisation capabilities, will serve as a pioneer on mm–wavelengths searches for pulsars and to study complex environments involving radio transients. To cover the northern sky, telescopes such as the IRAM 30m or the NOEMA observatory could be used, and in a future, the ngVLA.

### v4

@@ 4.2 Energy distribution
We constructed the cumulative energy distribution by sorting the events and then counting the number above an energy threshold, until all the pulses were contained. We included only events above the completeness limit 666This is the energy limit of a burst having the minimum detectable fluence given a telescope sensitivity., which is 1120 mJy ms using a 51 K and = 1.15 K Jy-1 (Liu et al. 2021) and considering the pulse width as the mean FWHM of our pulses, 0.07 s. We fit a power law to the cumulative energy distribution using a least-squares minimisation considering that the event rate scales with the energy as . The result is a power-law exponent of .
@@ 5 Discussion
Our sample adds five pulses in addition the three previously reported by Liu et al. (2021), mainly due to the difference in the candidate selection criteria. In our work, we defined the S/N threshold as 6, based on a statistical analysis of the noise floor. For each of these candidates the four Stokes parameters plus the linear and circular polarisation were inspected. Polarisation information was crucial as the dispersion delay between the frequency channels was negligible at ALMA Band 3 frequencies.
@@
Among the pulses, five of them (P1, P2, P3, P5, and P8) exhibit nearly 100 of the linearly polarised flux, while the remaining three (P4, P6, and P7) have values close to 50. In addition, for the majority of them there is a degree of circular polarisation of approximately 30. Regarding the pulse morphology, as seen in Fig. 1, most of the pulses have at least two components, which vary in relative intensity. All pulses show emission across the entire 2 GHz bandwidth. Broadband emission is one of the features of radio pulsars (Lorimer & Kramer 2004).
@@
This is the first time the energy properties of the pulses from the GC magnetar have been explored at high frequencies. Previous studies were conducted at 3.1 GHz (Yan et al. 2018) and 8.6 GHz (Yan et al. 2015). Common fits to the energy distribution of single pulses are log-normal and power-law distributions. For the power-law method, which is the one used in our study, the defining parameter is , the power-law exponent such that , with the rate of events above a given energy value . This definition corresponds to the cumulative energy distribution, which is broadly used in cases where the number of pulses is limited, such as the case of FRBs.
@@
As shown in Fig. 3, our sample is well fit by = -2.4 0.1. For the same source, at cm-wavelengths, the energy distribution of single pulses was found to be well fit by a log-normal distribution with = 1.34 and = 0.57 at 3.1 GHz (Yan et al. 2015), and = 0.72 and = 0.22 at 8.6 GHz (Yan et al. 2018). The difference with our result was expected given the sample size, where the authors considered 1900+ and 2300+ pulses, respectively. However, it is known that pulsars show a bimodal behaviour: their regular emission is well modelled with a log-normal distribution (Burke-Spolaor et al. 2012), while giant pulses are modelled by a power-law distribution (Karuppusamy et al. 2010). In the case of the GC magnetar, previous studies on the single pulse peak flux densities of giant pulses at 8.4 GHz, found that they are not well fit by a log-normal distribution (Pearlman et al. 2018). Given the steep spectral index in pulsars, the observable events at millimetre wavelengths could correspond to the most energetic pulses. Alternatively, a power-law relation for the cumulative energy distribution would be the result of the limited energy range of our pulses. Extending the observations with on-source time and frequency coverage will increase the sample and will provide us with a better general picture of the behaviour of PSR 1745-2900.
@@
Our is consistent with the energy distributions of other magnetar at cm-wavelengths. For XTE J1810197 (PSR J18091943) giant pulses, values between -2.1 and -7.68 at 1.4 GHz, between -1.78 and -2.86 at 4.9 GHz, and between -0.89 and -1.89 at 8.35 GHz have been reported. The differences could be attributed to its intrinsic magnetar variability and the observations at different frequencies (see Table 2 in Serylak et al. 2009).
@@
Importantly, our results fill a critical gap by characterising magnetar radio emission across a scarcely explored frequency window, offering valuable insights into potential changes in their emission mechanisms. Magnetar radio emission–known to be coherent at cm-wavelengths–persists into the mm-band, as demonstrated by our work, by Liu et al. (2021), and by Torne et al. (2017). However, it remains unclear exactly where a transition to an incoherent emission occurs. This transition is expected to take place in the mm to sub‑mm range; however, there are currently very few observational facilities equipped for time-domain astronomy at such frequencies. In this context, and given its sensitivity, PAM holds significant potential.
5.1 Comparison with FRBs
@@
In regard to FRBs, there are three sources with seemingly periodic activity windows: FRB 20121102A with a periodicity of 159.3 days (Braga et al. 2025), FRB 20180916B with 16.35 days (Chime/Frb Collaboration et al. 2020), and the recently reported FRB 20240209A with 126 days (Pal 2025). The periodicity corresponds to the cycle of the alternating on-and-off activity. Naturally, observations carried out at their peak activity are more likely to result in a detection. FRB 20121102A has values between (Cruces et al. 2021) and (Gourdji et al. 2019). However, Li et al. (2021) used a large sample of 1600+ bursts detected by the Five-hundred-metre Aperture Spherical Telescope (FAST), and showed that the energy distribution is bimodal and best fit by a combination of a log-normal and a Cauchy function. In that work, the analysis was done over the histogram of the differential energy distribution rather than the cumulative distribution, given the large sample of bursts available. In Li et al. (2021), when analysing the high-energy component, it was found that the sub-sample is well fit by a power law with . It is important to note that in our work we reported a limited sample of eight pulses. With a larger sample size (hundreds of pulses) the differential energy distribution could also be fit.
@@
For another active repeater, FRB 20201124A, -values of -1.5 0.1 (in Xu et al. 2022) and from -1.95 0.01 to -2.25 0.02 (in Kirsten et al. 2024) have been reported, both at cm-wavelengths. A steeper value of is often seen at higher energies in repeating FRBs (Kirsten et al. 2024), reaching values as steep as in the high-energy tail (Kumar et al. 2023). Recent studies of the FRB energy distribution have reported a value of -1.96 0.15 for the pure power law (Arcus et al. 2025) using FRB data between 0.83 and 1.6 GHz.
@@
The typical isotropic energies of FRBs are in the range of erg (CHIME/FRB Collaboration et al. 2020). Comparing the energy of our pulses, we find a gap of at least seven orders of magnitude. Nevertheless, the hypothesis of FRBs being extragalactic magnetars (Sobacchi et al. 2024) is supported by the activity of the Galactic magnetar SGR 1935+2154, which emitted bursts with energies in the range of 1034 (Margalit et al. 2019; Bochenek et al. 2020). This values are closer to those seen for FRBs, but still at least two orders of magnitude below. There are some phenomena that could explain the energy gap. One is multiple magnifications of the pulses due to plasma lensing, especially in environments where magnetars are orbiting a compact source such as a massive black hole (Pen & Connor 2015; Pearlman et al. 2018). In this regard, a promising laboratory to test this scenario is the GC. While for most radio pulsars and FRBs the monitoring is done at lower frequency bands, complex regions such as the GC are hard to reach with cm-waves as the pulses are heavily affected by propagation effects such as scattering and dispersion. Therefore higher frequencies allow us to probe deeper into these complex regions. The natural disadvantage is the lower luminosities expected at higher frequencies. Our study is especially remarkable given the detection of several single pulses from the GC magnetar at 86 GHz, which allow us to characterise the source, explore the potential of ALMA for radio transients and also to establish a direct link to what we could expect for FRBs at higher frequencies, if they are indeed originated by magnetars.
@@
Although most pulsars display a steep spectral index, magnetars such as PSR 1745-2900, have a flat or inverted one (Torne et al. 2017; Pearlman et al. 2018). Their spectral index could also change in time (see e.g. Champion et al. 2020). Torne et al. (2015) reports a value of for the GC magnetar, which changes to in a different epoch (Torne et al. 2017). This positive value could reflect a turn up, related with the transition from coherent to incoherent radio emission. Observing with high-sensitivity instruments such as ALMA could be helpful to study the gap between centimetre and millimetre radio wavelengths, particularly with its new Band 1 set-up.
5.2 PAM Band 1 predictions
@@
Based on our detections, we report a Poissonian rate of = 4 per hour (events/hr) for PSR J1745-2900 at ALMA Band 3. We can estimate the expected rate at Band 1 (42 GHz) using
@@
where corresponds to the expected event rate for a given fluence threshold at Band X. The fluence completeness limit using 37 antennas of 12m at Band 3 is = 1330 mJy ms and at Band 1 = 1200 mJy ms, considering a pulse width of 0.1 s. The exponent is the power-law value derived from the cumulative energy distribution discussed in Sect. 4.2. We obtain a rate of events/hr. However, if we instead consider a total of 49 antennas for the phasing, then the expected rate rises to events/hr, as the fluence threshold drops to = 900 mJy ms. Certainly, having a larger sample of pulses will provide more constraining values for their energy and frequency behaviour. In particular, measuring the spectral index of the source at different frequencies will tell us whether there is a turn-up point when we transition to incoherent emission mechanisms.
@@ 5.2.1 FRBs
If we consider the obtained for the GC magnetar and apply it to some actively repeating FRBs, we can estimate the event rates at ALMA bands. In this case, we re-calculated the fluence threshold for ALMA assuming bursts of 1 ms duration (typical for FRBs), and obtain = 90 mJy ms 777Using a maximum of 49 antennas of the same 12m size.. We take FRB 20121102A, whose reported rates at L-band are 122 events/hr (Li et al. 2021, 2022) and 21816 events/hr (Jahns et al. 2023). This leads to a rate between 1 1 and 3 2 events/hr, respectively. If on the contrary, we take the flatter value of reported for FRB 20121102A at L-band in Cruces et al. (2021), we estimate instead between 10 3 and 32 6 events/hr. It is noteworthy that these rates consider observations at the peak activity of the source.
We highlight that PAM Band 1 could serve as a bridge to millimetre wavelengths, which would have profound implications in constraining the FRBs’ progenitor source. If detections are achieved, it will also be key to test the proposed frequency dependence of the activity window (chromaticity) of periodic FRBs. As discussed in Bethapudi et al. (2023), the activity window seems to starts earlier and be narrower at higher frequencies. As shown in Fig. 4, currently ALMA could potentially detect 19 repeating FRBs. This number is expected to increase rapidly given the new surveying instruments in the southern hemisphere such as MeerKAT (Jonas & MeerKAT Team 2016).
@@ 5.2.2 Pulsars
We can also estimate the number of radio pulsars detectable in integrated pulse profile mode (also known as folding mode), which is obtained by adding hundreds to thousands of pulses using the spin period and other parameters listed in their ephemeris (see e.g. Table 3). We derived the minimum sensitivity for Band 3 using the radiometer equation for pulsars and considering telescope parameters such as = 74 K, = 1.08 K/Jy (Cortes et al. 2024) for Band 3, a phasing efficiency of (Matthews et al. 2018), a minimum detected S/N of 6 and 1 hour of on-source integration time. For a pulsar spin period of 1 s with a duty cycle of 10%, we estimate that the Band 3 system is sensitive to 0.060 mJy. If instead we use the parameters reported by Liu et al. (2021), we obtain 0.023 mJy. Assuming that the Band 1 system behaves similarly ( = 45 K, = 1.12 K/Jy in Cortes et al. 2024), then we expect a minimum detectable flux of 0.035 mJy, using the same 37-antenna set-up. When increasing to 49 antennas, the minimum flux is 0.026 mJy.
In Fig. 5 we show the sources expected to be detectable by PAM Band 1. To this end, we selected all the known pulsars from the ATNF catalogue (Manchester et al. 2005) having flux density reported at 1.4 GHz or 2 GHz. We limited the sources detectable by ALMA based on a declination below 47° (Cortes et al. 2024). This reduced the list to 2301 pulsars.
We used the spectral index to estimate the flux of each source at 42 GHz (Band 1). To extrapolate we used a simple power-law:
Here corresponds to the flux at a given frequency in GHz, is the spectral index, and is the frequency. For the sources that do not have a reported spectral index, we draw a spectral index from a normal distribution, considering a mean of -1.6 and standard deviation equal to 0.54 (Jankowski et al. 2017b). We assumed a pulse width equal to 10 of the period and considered on-source integration times of 1, 1.5, and 2.5 hours. This would lead to the detection of 166 pulsars (7.2 of those visible to ALMA), 188 (8.2), and 209 (9.1), respectively. The sources detectable by PAM are the ones lying above the mentioned sensitivity lines in Fig. 5. Some remarkable detectable sources have been highlighted, such as the Vela pulsar, and three radio-loud magnetars, including PSR J17452900. It is important to mention that these surveys are unlikely to be sensitive to millisecond pulsars (MSPs, Liu et al. 2021).
@@ 6 Conclusions
We reprocessed archival data of APS at Band 3 from Sgr A∗. We detected eight pulses from the GC magnetar PSR J17452900 and studied their cumulative energy distribution with the objective to link this with FRBs. The complex environment where the GC magnetar is located serves us to test promising progenitor scenarios for FRBs, given its high electron density and turbulent magneto ionic properties.
We obtained an exponent of = 2.4 0.1 for the cumulative energy distribution of the GC magnetar. Using this exponent we predict a rate of 5 events/hr if we were to observe in Band 1, using 37 antennas of 12m. This rate increases to 10 events/hr with 49 12m antennas. For the periodic FRB 20121102A, at peak activity, we obtained a rate between 1 and 32 events/hr with PAM Band 1 using 49 12m antennas. Regarding pulsar detection, we estimated that there are roughly 166 pulsars potentially detectable within 1 hr integration. This number increases with more integration time.
Currently, there are 19 repeating FRBs that are visible to ALMA. Given its high sensitivity and time resolution, instruments like PAM are crucial to exploring the FRB frequency extent to test their progenitor scenarios. Importantly, high-frequency searches are not affected by propagation effects, such as scattering, which could hinder a detection at cm-wavelengths. It is expected that the Band 1 receiver, including polarisation capabilities, will serve as a pioneer on mm-wavelengths searches for pulsars and to study complex environments involving radio transients. To cover the northern sky, telescopes such as the IRAM 30m or the NOEMA observatory could be used and, in the future, the ngVLA.